# Congestion-Aware Path Re-routing Strategy for Dense Urban Airspace

### v1

@@ IV-A2 Markov modulated Bernoulli process (MMBP)
IV-A3 Queue overflow
The UAS that have queued for $L+1$ slots are lost from the queue and become non-nodal arrivals $a_{\mathrm{I},k}$ to the Stream ±$\pm 1$ queueing system. The steady-state probability that in timeslot $k$, a UAS queuing would be lost from the system is called Queue overflow probability $\Phi$. Let $q_{k}\in\{0,1\}$ be a random variable that there exists a UAS that has queued for $L$ slots in time slot $k$.
@@ IV-A3 Queue overflow
$\mathrm{P}[v_{L,k}=j]\ =\ \mathrm{P}[v_{L-1,k}=j-1]\ \mathrm{P}[q_{k}=1]\ +$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \mathrm{P}[v_{L-1,k}=j]\ \mathrm{P}[q_{k}=0]\ ,\ j=1,...,L-1$ (29)
@@
Summing Eqns. 29-30 for $j=1,...,L$ and simplifying we get,
$\sum_{j=1}^{L}\mathrm{P}[v_{L,k}=j]\ =\ \mathrm{P}[q_{k}=1]\ \sum_{j=1}^{L}\mathrm{P}[v_{L-1,k}=j-1]$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\mathrm{P}[q_{k}=0]\ \sum_{j=1}^{L-1}\mathrm{P}[v_{L-1,k}=j-1]$ (31)
@@
The UAS corresponding to $q_{k}=1$ has the least preference ($\psi_{Z}=-L$) amongst other UAS present in the queue, including the nodal arrival $(\psi_{Z}=0,-1,...,-L+1)$. If the $Z_{\mathrm{UP}}$ is uncongested in time slot $k$ and there are higher $\psi_{Z}$ UAS, then the UAS that has last arrived would enter service, forcing the UAS corresponding to $q_{k}=1$ to queue an additional slot. This UAS would have queued for $L+1$ slots and hence is lost from the system. When $Z_{\mathrm{UP}}$ turns congested, under steady-state, the fraction of UAS arrivals that entered the queue but failed to enter service when $Z_{\mathrm{UP}}$ was uncongested are lost during the congestion period. Thus, the Queue overflow probability is given by,
$\Phi$ $=(1-\Theta_{0}^{*})\left[1-A(0)+A(0)(1-V_{L-1}(0))\right]\ \mathrm{P}[q_{k}=1]$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ +\Theta_{0}^{*}\ (1-A(0))(1-V_{L}(0))W_{1}^{*}(0)$ (34)
Following Rule 7, the steady-state probability distribution of the non-nodal arrivals $a_{\mathrm{I},k}$ for the neighboring Stream $-1$ queuing system is given by
@@
The non-nodal arrival probability distribution $A_{\mathrm{I}}(z)$ for neighboring Stream $+1$ system is given by
@@
IV-B Stream $X$ queuing system
The schematic for the Stream $X$ queuing system is shown in Fig. 17. A Stream $X$ queuing system $\overline{Z}$ has nodal arrivals $\{a_{k}\}$ and non-nodal arrivals $\{a_{\mathrm{I},k},a_{\mathrm{O},k}\}$. These arrivals are departures and overflows from the preceding systems. Assume that the PGF of these arrivals has already been computed. Let $b_{k}$ be a random variable such that $b_{k}=1$ if a nodal arrival in time slot $k$ would descend to $Z_{\mathrm{UD}}$ zone and is 00 otherwise. Corresponding descend probability $\sigma_{\mathrm{I}}$ and PGF $B(z)=(1-\sigma_{\mathrm{I}})+z\sigma_{\mathrm{I}}$ are also known. $B(z)$ is dependent on the arrivals and UAS queuing in the $Z_{\mathrm{IN}}$ system (refer Eqn. 21). The non-nodal arrival $a_{\mathrm{I},k}$ on finding the $\overline{Z}$ system busy would enter the $\gamma_{Z}$ queue. The nodal arrival $a_{k}$ on finding the $\overline{Z}$ system busy and descend condition unsatisfied would enter the $\beta_{Z}$ queue.
Figure 17: Stream $X\neq 0$ queuing system
Let ${}^{\gamma}v_{j,k}\in\{0,1,...,j\},j\in\{1,...,L-1\}$ be random variables such that at time slot $k$, ${}^{\gamma}v_{j,k}$ gives the total number of UAS currently in the $\gamma_{Z}$ queue that have queued for no more than $j$ slots since entering the queue. ${}^{\gamma}V_{j}(z)$ are respective PGFs. ${}^{\gamma}v_{L-1,k}$ is the total number of UAS queuing in the $\gamma_{Z}$ queue. The non-nodal arrival $a_{\mathrm{I},k}$ on finding $Z_{\mathrm{UP}}$ zone congested for $L$ successive slots would enter the node $\widetilde{Z}$. Thus, for UAS queuing in the $\gamma_{Z}$ queue, there is a $L-1$ slot queuing deadline to enter service, beyond which these UAS are lost from the $\gamma_{Z}$ queue. The UAS that are lost from the $\gamma_{Z}$ queue re-enter the system via node $\widetilde{Z}$. Hence, these UAS are also treated as nodal arrivals $c_{k}$ to the system with PGF $C(z)$. If there are simultaneous $c_{k}$ and $a_{k}$ nodal arrivals in timeslot $k$, then there is conflict at node $\widetilde{Z}$. In conflict, the nodal arrival $c_{k}$ has the highest preference and enters service. Whereas the non-nodal arrival $a_{\mathrm{I},k}$ and the UAS queuing in $\gamma_{Z}$ queue are forced to queue an additional slot. In timeslots with no conflict, the UAS in the $\gamma_{Z}$ queue that has queued for the least number of slots has the highest preference and enters service. This is nothing but the LCFS discipline.
Let ${}^{\gamma}q_{k}\in\{0,1\}$ be a random variable that there exists a UAS that has queued for $L-1$ slots in $\gamma_{Z}$ queue in time slot $k$. When $Z_{\mathrm{UP}}$ is uncongested, the UAS corresponding to $q_{k}=1$ is lost from the system if,
@@ IV-B Stream $X$ queuing system
    there is a conflict at node $\widetilde{Z}$ or
@@
    there are no conflicts at node $\widetilde{Z}$, but there are non-nodals arrivals $a_{\mathrm{I},k}=1$ or UAS in $\gamma_{Z}$ queue (${}^{\gamma}v_{L-2,k}>0$) that have higher preference than UAS $q_{k}=1$.
When $Z_{\mathrm{UP}}$ turns congested, then under steady-state, the fraction of UAS arrivals that entered the $\gamma_{Z}$ queue but failed to enter service when $Z_{\mathrm{UP}}$ was uncongested are lost during the congestion period. The above behavior is similar to that of Stream 00 queue. Thus, similar to Eqns. 33-34, the PGF of nodal arrivals $c_{k},k\longrightarrow\infty$ is given as follows.
${}^{\gamma}v_{L-1,k}$ $={}^{\gamma}v_{L-2,k}+q_{k}$ (37)
$\mathrm{P}[q_{k}=1]$ $=\left(\cfrac{{}^{\gamma}V_{L-2}(0)-{}^{\gamma}V_{L-1}(0)}{{}^{\gamma}V_{L-2}(0)}\right)$ (38)
$\Omega_{0}$ $=(1-A(0))\ \mathrm{P}[q_{k}=1]$ (39)
where $\Omega_{0}$ is a pessimistic probability of conflict in timeslot $k$.
$\mathrm{P}[c_{k}=1]$ $=(1-\Theta_{0})\left[\ \Omega_{0}+(1-\Omega_{0})(\ (1-A_{\mathrm{I}}(0))\right.$
$\left.+\ A_{\mathrm{I}}(0)(1-{}^{\gamma}V_{L-2}(0))\ \right]\ \mathrm{P}[q_{k}=1]$
$+\Theta_{0}(1-A_{\mathrm{I}}(0))(1-{}^{\gamma}V_{L-1}(0))W_{1}(0)$ (40)
@@
The LCFS dynamic equations for UAS queuing in the $\gamma_{Z}$ queue are as follows. If the system is busy ($\theta_{k}=0$) or there is a conflict ($a_{k}=c_{k}=1$) then
${}^{\gamma}v_{1,k+1}=a_{\mathrm{I},k}$ (42)
${}^{\gamma}v_{j+1,k+1}={}^{\gamma}v_{j,k}+a_{\mathrm{I},k},\ j\in\{1,...,L-1\}$ (43)
@@
${}^{\gamma}v_{1,k+1}=(a_{\mathrm{I},k}-1)^{+}=0$ (44)
${}^{\gamma}v_{j+1,k+1}=({}^{\gamma}v_{j,k+1}+a_{\mathrm{I},k}-1)^{+},\ j\in\{1,...,L-2\}$ (45)
Eqns. 42-45 are similar to Eqns. 14-17. Thus, ${}^{\gamma}V_{j}(z),\ j\in\{1,...,L-1\}$ for the $\gamma$ queue can be derived using the conditional PGF as,
${}^{\gamma}V_{1}(z)=A_{\mathrm{I}}(z)(\Theta_{0}+\Omega_{0}-\Theta_{0}\Omega_{0})$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +(1-\Theta_{0})(1-\Omega_{0})$ (46)
${}^{\gamma}V_{j+1}(z)=A_{\mathrm{I}}(z){}^{\gamma}V_{j}(z)(\Theta_{0}+\Omega_{0}-\Theta_{0}\Omega_{0})$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ +E[z^{({}^{\gamma}v_{j,k}+a_{\mathrm{I},k}-1)^{+}}](1-\Theta_{0})(1-\Omega_{0})$ (47)
By recursively substituting ${}^{\gamma}V_{j}(z)$ in ${}^{\gamma}V_{j+1}(z)$, we get the PGF ${}^{\gamma}V_{L-1}(z)$. The nodal arrivals $c_{k}$ and $a_{k}$ that fail to enter service in time slot $k$ would enter the $\beta_{Z}$ queue in the next time slot. Let ${}^{\beta}v_{j,k}\in\{0,1,...,j\},\ j\in\{1,...,L\}$ be random variables such that at time slot $k$, ${}^{\beta}v_{j,k}$ give the total number of UAS currently in the $\beta_{Z}$ queue, that have queued for no more than $j$ slots since entering the queue. ${}^{\beta}V_{j}(z)$ are respective PGFs. ${}^{\beta}v_{L,k}$ is the total number of UAS queuing in the $\beta_{Z}$ queue. The queuing behavior of the $\beta_{Z}$ queue when the system is busy or idle, when the descend condition is satisfied or not, and when there is a conflict at node $\widetilde{Z}$ or not can be summarized in the following eight scenarios. Note that $a_{k},\ c_{k}$ and ${}^{\beta}v_{j,k}$ are independent at timeslot $k$. For $j\in\{1,...,L-1\}$,
@@
    When there is a conflict ($a_{k}=c_{k}=1$) due to simultaneous $a_{k}$ and $c_{k}$ arrivals and the descend condition is satisfied ($b_{k}=1$) with probability $(1-C(0))(1-A(0))(1-B(0))$, then by rule 5 the $c_{k}$ arrival enters service and by rule 6 the $a_{k}$ descends to $Z_{\mathrm{UD}}$ zone. No arrival enters the $\beta_{Z}$ queue. Irrespective of whether the system is busy or not and whether UAS are present in the $\gamma_{Z}$ queue or not, the UAS present in the $\beta_{Z}$ queue are forced to queue an additional slot.
${}^{\beta}v_{1,k+1}=0\ ,$ ${}^{\beta}V_{1}(z)=1$ (48)
${}^{\beta}v_{j+1,k+1}={}^{\beta}v_{j,k}\ ,$ ${}^{\beta}V_{j+1}(z)={}^{\beta}V_{j}(z)$ (49)
@@
    If there is a conflict ($a_{k}=c_{k}=1$) but the descend condition is not satisfied ($b_{k}=0$) with probability $(1-C(0))(1-A(0))B(0)$, then the arrival $c_{k}$ would enter service and $a_{k}$ would enter $\beta_{Z}$ queue.
$\ \ \ {}^{\beta}v_{1,k}=1\ ,$ ${}^{\beta}V_{1}(z)=z$ (50)
$\ \ \ {}^{\beta}v_{j+1,k+1}={}^{\beta}v_{j,k}+1\ ,$ ${}^{\beta}V_{j+1}(z)=z\ {}^{\beta}V_{j}(z)$ (51)
  1. 3.
    If either the system is busy ($\theta_{k}=0$) or idle ($\theta_{k}=1$) with UAS having higher preference being present ($a_{\mathrm{I},k}=1$ or ${}^{\gamma}v_{L-1,k}>0$) with probability $\rho$ then the UAS present in $\beta_{Z}$ queue do not enter service.
    $\ \ \rho\ =$ ${[}\ \Theta_{0}+(1-\Theta_{0}){(}(1-A_{\mathrm{I}}(0))$
    $\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +A_{\mathrm{I}}(0)(1-{}^{\gamma}V_{L-1}(0)){)}\ {]}$ (52)
@@
      If there is no conflict because $a_{k}=0$ and $c_{k}=\{0,1\}$ with probability $\rho A(0)$, then the arrival $c_{k}$ would enter the $\beta_{Z}$ queue.
      ${}^{\beta}v_{1,k}=c_{k}\ ,$ ${}^{\beta}V_{1}(z)=C(z)$ (53)
      ${}^{\beta}v_{j+1,k+1}={}^{\beta}v_{j,k}+c_{k}\ ,$ ${}^{\beta}V_{j+1}(z)={}^{\beta}V_{j}(z)C(z)$ (54)
@@
      If there is no conflict because $c_{k}=0$, $a_{k}=\{0,1\}$ and
@@
        if descend condition is satisfied ($b_{k}=1$) with probability $\rho C(0)(1-B(0))$ then no UAS will enter the $\beta_{Z}$ queue.
        ${}^{\beta}v_{1,k}=0\ ,$ ${}^{\beta}V_{1}(z)=1$ (55)
        ${}^{\beta}v_{j+1,k+1}={}^{\beta}v_{j,k}\ ,$ ${}^{\beta}V_{j+1}(z)={}^{\beta}V_{j}(z)$ (56)

### v2

@@ 4.1.2 Markov modulated Bernoulli process (MMBP)
4.1.3 Queue overflow
The UAS that have queued for $L+1$ slots are lost from the queue and become non-nodal arrivals $a_{\mathrm{I},k}$ to the neighboring Stream{±$\{\pm 1\}$ queueing systems. The steady-state probability that in timeslot $k$, a UAS queuing would be lost from the system is called Queue overflow probability $\Phi$. Let $q_{k}\in\{0,1\}$ be a random variable such that in the given timeslot $k$, $q_{k}=1$ denotes that there exists a UAS that has queued $L$ timeslots and 00 otherwise.
@@ 4.1.3 Queue overflow
$\mathrm{P}[v_{L,k}=j]\ =\ \mathrm{P}[v_{L-1,k}=j-1]\ \mathrm{P}[q_{k}=1]\ +\mathrm{P}[v_{L-1,k}=j]\ \mathrm{P}[q_{k}=0]\ \ \ ,\ \ \ j=1,...,L-1$ (23)
@@
Summing Eqns. 23-24 for $j=1,...,L$ and simplifying we get,
$\sum_{j=1}^{L}\mathrm{P}[v_{L,k}=j]\ =\ \mathrm{P}[q_{k}=1]\ \sum_{j=1}^{L}\mathrm{P}[v_{L-1,k}=j-1]+\mathrm{P}[q_{k}=0]\ \sum_{j=1}^{L-1}\mathrm{P}[v_{L-1,k}=j-1]$ (25)
@@
The UAS corresponding to $q_{k}=1$ has the least preference ($\psi_{Z}=-L$) amongst other UAS present in the queue $(\psi_{Z}\in\{0,-1,...,-L+1\})$, including the nodal arrival. If the $Z_{\mathrm{UP}}$ is uncongested in timeslot $k$ and there is at least one UAS in the queue that has preference $\psi_{Z}>-L$. Then, by priority, the UAS that has the higher preference enters service, forcing the UAS corresponding to $q_{k}=1$ to queue in the given timeslot. This UAS would have queued for $L+1$ slots and hence is lost from the system. In steady-state, when $Z_{\mathrm{UP}}$ is congested, the fraction of UAS arrivals that entered the queue but failed to enter service when $Z_{\mathrm{UP}}$ was uncongested are lost during the congestion period. Notice that when the $Z_{\mathrm{UP}}$ is congested, as well as when uncongested, the UAS may exit the system. Thus, the Queue overflow probability is given by,
[…]
The queue overflows and departures are correlated processes that are better modeled as a higher-order Markov model (order $L$ or higher) with covariate dependencies in transition probabilities (Ataharul Islam and Chowdhury (2006)). However, for ease of analysis, we use a first-order Markov model and compensate for modeling errors by introducing an empirical correction factor $\xi(x)$ as in Eqn. 28. The factor $\xi(x)$ is significant only when the UAS arrival rate $x$ is low.
$\xi(x)=0.15\left(\cfrac{x}{M-1+\zeta}\right)\exp\left({-S\left(\cfrac{x}{M-1+\zeta}\right)^{3}}\right)\approx(2-10\%)x\ \ ,\ \ \zeta=\begin{cases}2\eta&\text{if\ }\eta\in[0,0.5]\\ 2(1-\eta)&\text{if\ }\eta\in[0.5,1]\end{cases}$ (29)
Following Rule 8, the steady-state probability distribution of the non-nodal arrivals $a_{\mathrm{I},k}$ for the neighboring Stream$\{-1\}$ queuing system is given by
@@
The non-nodal arrival probability distribution $A_{\mathrm{I}}(z)$ for neighboring Stream$\{+1\}$ system is given by
@@
4.2 Stream$\{X\}$ queuing system, $X\neq 0$
The schematic for the Stream$\{X\}$ queuing system associated with Stream$(X)$ zone is shown in Fig. 14. A Stream$\{X\}$ queuing system $\overline{Z}$ has nodal arrivals $\{a_{k}\}$ and non-nodal arrivals $\{a_{\mathrm{I},k},a_{\mathrm{O},k}\}$. These arrivals are departures and overflows from the preceding systems. Assume that the PGF of these arrivals have already been computed. Let $b_{k}$ be a random variable such that $b_{k}=1$ if a nodal arrival in timeslot $k$ would descend to $Z_{\mathrm{ID}}$ zone and is 00 otherwise. Corresponding descend probability $\sigma_{\mathrm{I}}$ and PGF $B(z)=(1-\sigma_{\mathrm{I}})+z\sigma_{\mathrm{I}}$ are available, as $\sigma_{\mathrm{I}}$ is dependent only on the arrivals and UAS queuing in the $\overline{Z}_{\mathrm{IN}}$ system (refer Eqn. 16). The non-nodal arrival $a_{\mathrm{I},k}$, on finding the $\overline{Z}$ system busy, would enter the $\beta_{Z}$ queue. The nodal arrival $a_{k}$, on finding the $\overline{Z}$ system busy and descend condition unsatisfied, would enter the $\gamma_{Z}$ queue.
Figure 14: Stream$\{X\}$ queuing system, $X\neq 0$
Let ${}^{\beta}v_{j,k}\in\{0,1,...,j\},\ j\in\{1,...,L-1\}$ be random variables such that at timeslot $k$, ${}^{\beta}v_{j,k}$ gives the total number of UAS currently in the $\beta_{Z}$ queue that have queued for no more than $j$ slots since entering the queue. Here, ${}^{\beta}V_{j}(z)$ are respective PGFs and, ${}^{\beta}v_{L-1,k}$ is the total number of UAS queuing in the $\beta_{Z}$ queue. The non-nodal arrival $a_{\mathrm{I},k}$ on finding $Z_{\mathrm{UP}}$ zone congested for $L$ successive slots would enter the node $\widetilde{Z}$. Thus, for UAS queuing in the $\beta_{Z}$ queue, there is a $L-1$ slot queuing deadline to enter service, beyond which these UAS are lost from the $\beta_{Z}$ queue. The UAS that exits the $\beta_{Z}$ queue re-enters the system via node $\widetilde{Z}$. Hence, these UAS are also treated as nodal arrivals $c_{k}$ to the system with PGF $C(z)$. If there are simultaneous $c_{k}$ and $a_{k}$ nodal arrivals in timeslot $k$, then there is conflict at node $\widetilde{Z}$. In conflict, the nodal arrival $c_{k}$ enters service (Rule 6). Whereas the non-nodal arrival $a_{\mathrm{I},k}$ and the UAS queuing in $\beta_{Z}$ queue in the given timeslot. In timeslots with no conflict, the UAS in the $\beta_{Z}$ queue that has queued for the least number of slots has the highest preference and enters service. This is the LCFS discipline.
Let ${}^{\beta}q_{k}\in\{0,1\}$ be a random variable that denotes if there exists a UAS that has queued for $L-1$ slots in $\beta_{Z}$ queue in timeslot $k$. When $Z_{\mathrm{UP}}$ is uncongested, the UAS corresponding to ${}^{\beta}q_{k}=1$ is lost from the $\beta_{Z}$ queue if,
@@ 4.2 Stream$\{X\}$ queuing system, $X\neq 0$
    there are non-nodal arrivals ($a_{\mathrm{I},k}=1$) or
@@
    there are UAS in $\beta_{Z}$ queue (${}^{\beta}v_{L-2,k}>0$) that have higher preference than the UAS corresponding to ${}^{\beta}q_{k}=1$.
In steady-state, when $Z_{\mathrm{UP}}$ is congested, the fraction of UAS arrivals that entered the $\beta_{Z}$ queue but failed to enter service when $Z_{\mathrm{UP}}$ was uncongested are lost during the congestion period. The above behavior is similar to that of Stream$\{0\}$ queue, where $c_{k}$ arrivals are queue overflows from $\beta_{Z}$ queue. Thus, similar to Eqns. 27-28, the PGF of nodal arrivals $c_{k},\ k\longrightarrow\infty$ is given as follows.
${}^{\beta}v_{L-1,k}$ $=\ {}^{\beta}v_{L-2,k}+{}^{\beta}q_{k}\ \ \ \ ,\ \ \ \ \mathrm{P}\left[{}^{\beta}q_{k}=1\right]=\left(\cfrac{{}^{\beta}V_{L-2}(0)-{}^{\beta}V_{L-1}(0)}{{}^{\beta}V_{L-2}(0)}\right)$ (32)
$\mathrm{P}[c_{k}=1]$ $=(1-\Theta_{0})\left[1-A_{\mathrm{I}}(0){}^{\beta}V_{L-2}(0)\right]\mathrm{P}\left[{}^{\beta}q_{k}=1\right]+\Theta_{0}(1-A_{\mathrm{I}}(0))\left(1-{}^{\beta}V_{L-1}(0)\right)W_{1}(0)+\xi(1-A_{\mathrm{I}}(0))$ (33)
@@
The LCFS dynamic equations for UAS queuing in the $\beta_{Z}$ queue are as follows. If the system is busy ($\theta_{k}=0$) or there is a conflict ($a_{k}=c_{k}=1$) then
${}^{\beta}v_{1,k+1}=a_{\mathrm{I},k}\ \ \ \ ,\ \ \ \ {}^{\beta}v_{j+1,k+1}={}^{\beta}v_{j,k}+a_{\mathrm{I},k}\ \ \ \ ,\ \ \ \ j\in\{1,...,L-2\}$ (35)
@@
${}^{\beta}v_{1,k+1}=(a_{\mathrm{I},k}-1)^{+}=0\ \ \ \ ,\ \ \ \ {}^{\beta}v_{j+1,k+1}=({}^{\beta}v_{j,k+1}+a_{\mathrm{I},k}-1)^{+}\ \ \ \ ,\ \ \ \ \ j\in\{1,...,L-2\}$ (36)
Eqns. 35-36 are similar to Eqns. 12-13. Thus, ${}^{\beta}V_{j}(z),\ j\in\{1,...,L-1\}$ for the $\beta$ queue can be derived using the conditional PGF as,
$\Omega=(1-C(0))(1-A(0))$ (37)
${}^{\beta}V_{1}(z)=A_{\mathrm{I}}(z)(\Theta_{0}+\Omega-\Theta_{0}\Omega)+(1-\Theta_{0})(1-\Omega)$ (38)
${}^{\beta}V_{j+1}(z)=A_{\mathrm{I}}(z){}^{\beta}V_{j}(z)(\Theta_{0}+\Omega-\Theta_{0}\Omega)+\left[{}^{\beta}V_{j}(0)+\left(\cfrac{{}^{\beta}V_{j}(z)-{}^{\beta}V_{j}(0)}{z}\right)A_{\mathrm{I}}(z)\right](1-\Theta_{0})(1-\Omega)$ (39)
where $\Omega$ is the conflict probability. The conflict probability $\Omega$ is dependent on the probability $(1-C(0))$, which is not known. If $\Omega=0$, then Eqns. 38-39 are same as the Eqn. 15 of Stream$\{0\}$ queuing system. Assuming $\Omega=0$, by substituting $A(z)=A_{\mathrm{I}}(z)$ in Eqn. 15 and solving for queue overlow probability $\Phi$ using Eqns. 15-28, we get a good estimate for the probability $(1-C(0))=\Phi$. Then, conflict probability $\Omega$ can be computed using Eqn. 37. By recursively substituting ${}^{\beta}V_{j}(z)$ in ${}^{\beta}V_{j+1}(z)$, we get the PGF ${}^{\beta}V_{L-1}(z)$. The nodal arrivals $c_{k}$ and $a_{k}$ that fail to enter service in timeslot $k$ would enter the $\gamma_{Z}$ queue in the next timeslot. Let ${}^{\gamma}v_{j,k}\in\{0,1,...,j\},\ j\in\{1,...,L\}$ be random variables such that at timeslot $k$, ${}^{\gamma}v_{j,k}$ give the total number of UAS currently in the $\gamma_{Z}$ queue, that have queued for no more than $j$ timeslots since entering the queue. Here, ${}^{\gamma}V_{j}(z)$ are respective PGFs and, ${}^{\gamma}v_{L,k}$ is the total number of UAS queuing in the $\gamma_{Z}$ queue. The queuing behavior of the $\gamma_{Z}$ queue when the system is busy or idle, when the descend condition is satisfied or not, and when there is a conflict at node $\widetilde{Z}$ or not can be summarized in the following six scenarios. Note that $a_{k},\ c_{k}$ and ${}^{\gamma}v_{j,k}$ are independent at timeslot $k$. For $j\in\{1,...,L-1\}$,
@@
    If there is a conflict ($a_{k}=c_{k}=1$) due to simultaneous nodal arrivals $a_{k}$ and $c_{k}$ and,
    1. (a)
      the descend condition is satisfied ($b_{k}=1$) with probability $\Omega(1-B(0))$, then by Rule 6, the nodal arrival $c_{k}$ enters service and the nodal arrival $a_{k}$ descends to $Z_{\mathrm{ID}}$ zone. No arrival enters the $\gamma_{Z}$ queue. The UAS present in the $\gamma_{Z}$ queue would queue in the given timeslot.
      ${}^{\gamma}v_{1,k+1}=0\ \ ,\ \ {}^{\gamma}v_{j+1,k+1}={}^{\gamma}v_{j,k}\ \ \ \ \implies\ \ \ \ {}^{\gamma}V_{1}(z)=1\ \ ,\ \ {}^{\gamma}V_{j+1}(z)={}^{\gamma}V_{j}(z)$ (40)
    2. (b)
      the descend condition is not satisfied ($b_{k}=0$) with probability $\Omega B(0)$, then the arrival $c_{k}$ would enter service and arrival $a_{k}$ would enter $\gamma_{Z}$ queue.
      ${}^{\gamma}v_{1,k}=1\ \ ,\ \ {}^{\gamma}v_{j+1,k+1}={}^{\gamma}v_{j,k}+a_{k}\ \ \ \ \implies\ \ \ \ {}^{\gamma}V_{1}(z)=A(z)\ \ ,\ \ {}^{\gamma}V_{j+1}(z)=A(z)\ {}^{\gamma}V_{j}(z)$ (41)
@@
    If there is no conflict ($a_{k}=c_{k}=0$ or $a_{k}\neq c_{k}$), that is, both $a_{k}=1$ and $c_{k}=1$ do not occur simultaneously and,
@@
      if the system is busy ($\theta_{k}=0$) or if the system is idle ($\theta_{k}=1$) but there are UAS with higher preference present ($a_{\mathrm{I},k}=1$ or ${}^{\beta}v_{L-1,k}>0$) with probability $\rho$ then the UAS in the $\gamma_{Z}$ queue do not enter service and,
[…]
      1. i.
        if the descend condition is satisfied ($b_{k}=1$), then nodal arrival $a_{k}$ descends to $Z_{\mathrm{ID}}$ zone and the nodal arrival $c_{k}$ enters the $\gamma_{Z}$ queue with probability $\rho(1-B(0))(1-\Omega)$.
        ${}^{\gamma}v_{1,k}=c_{k}\ ,\ {}^{\gamma}v_{j+1,k+1}={}^{\gamma}v_{j,k}+c_{k}\ \ \implies\ \ {}^{\gamma}V_{1}(z)=C(z)\ ,\ {}^{\gamma}V_{j+1}(z)=\ {}^{\gamma}V_{j}(z)C(z)$ (43)
      2. ii.
        if the descend condition is not satisfied ($b_{k}=0$), then nodal arrivals $a_{k}$ and $c_{k}$ enter the $\gamma_{Z}$ queue with probability $\rho B(0)(1-\Omega)$.
        ${}^{\gamma}v_{1,k}=a_{k}+c_{k}$ $,\ \ \ \ {}^{\gamma}v_{j+1,k+1}=\ {}^{\gamma}v_{j,k}+a_{k}+c_{k}$
[…]
@@
      if the system is idle ($\theta_{k}=1$) and no UAS with higher preference is present ($a_{\mathrm{I},k}=0$ and ${}^{\beta}v_{L-1,k}<0$) with probability $(1-\rho)$ and,
@@
        if the descend condition is satisfied ($b_{k}=1$) with probability $(1-\rho)(1-B(0))(1-\Omega)$, then nodal arrival $a_{k}$ descends to $Z_{\mathrm{ID}}$. If nodal arrival $c_{k}=1$, then following LCFS discipline, the nodal arrival enters service. Otherwise, the UAS in the $\gamma$ queue that has arrived last enters service, thus reducing the number of UAS queuing in the $\gamma$ queue by one.
        ${}^{\gamma}v_{1,k}=(c_{k}-1)^{+}=0\ \ ,$ ${}^{\gamma}v_{j+1,k+1}=\ {}^{\gamma}(v_{j,k}+c_{k}-1)^{+}$
[…]